Special Text Characters

Type: Print: Type:  Print: Type: Print
\#  # \$  $ \%  %
\&  & \˜ {}  ˜  \
\ˆ {}  ˆ  \{  { \}  }
$|$  |  @  @ $∗$  ∗
$\backslash$ \

Text Accents

Type: Print: Type: Print: Type: Print
\‘{o} ò \´ {o} ó \”{o} ö
\H{o} ő \ˆ {o} ô \˜ {o}  õ
\v{o}  ǒ \u{o}  ŏ \={o} ō
\b{o}  ō \.{o}  ȯ \d{o}  o.
\c{o}  o̧ \r{o}  o̊ \t{oo} oo
\i  ı  \j  ȷ

Some European characters

Type: Print: Type: Print: Type: Print
\aa  å \AA  Å \ae  æ
\AE  Æ \o  ø \O  Ø
\oe  œ \OE  Œ \l  ł
\L  Ł \ss  ß \SS  SS
? ‘  ¿  ! ‘  ¡

Extra Text Symbol

Type:  Print:
\dag  †
\ddag  ‡
\S  §
\P  ¶
\copyright \begin{matrix}c \\ ◯ \end{matrix}
\pounds  £
\textbullet  •

\textvisiblespace

\textcircled\{a\} \begin{matrix}a \\ ◯ \end{matrix}
\textperiodcentered  ·

Text Spacing Commands

Short form: Full form:  Size: Short form: Full form:
\,  \thinspace  a a \!  \negthinspace
\:  \medspace  a a  \negmedspace
\;  \thickspace a a  \negthickspace
\quad  a a
\qquad  a a

The \medspace, \thickspace, \negmedspace, and \negthickspace commands require the amsmath package.

Text Font Commands

command  command  switch to
with argument  declaration
\textnormal{ … } {\normalfont … } document font family
\textrm{ … }  {\rmfamily … } roman font family
\textsf{ … }  {\sffamily … }  sans serif font family
\texttt{ … }  {\ttfamily … }  typewriter style font family
\textup{ … }  {\upshape … }  upright shape
\textit{ … }  {\itshape … }  italic shape
\textsl{ … }  {\slshape … }  slanted shape
\textnormal{ … } {\normalfont … } default font
\textsc{ … }  {\scshape … }  small capitals
\emph{ … }  {\em … }  emphasis
\textbf{ … }  {\bfseries … }  bold extended
\textmd{ … }  {\mdseries … }  normal weight and width

Text Font Sizes Changes

\tiny  sample text
\scriptsize  sample text
\footnotesize  sample text
\small  sample text
\normalsize  sample text
\large  sample text
\Large  sample text
\huge  sample text
\Huge sample text